OpenStudy (anonymous):

A motorboat can go 14 miles downstream on a river in 20 minutes. it takes 30 minutes for this boat to go back upstream the same 14 miles. Find the speed of the boat. Cheryl is three years more than three times as old as Ben. In 20 years the sum of their ages will be 87. How old is Ben now? An eastbound train leaves Vancouver for Toronto, a distance of about 4370 km, travelling at about 90km/h. At the same time a west bound train leaves Toronto to Vancouver travelling at about 70 km/h. How far from Toronto are the trains when they pass each other? For how long have they been travelling?

OpenStudy (perl):

for the first problem Use distance = rate * time Assume that r = speed of boat in still water w = water current 14 = ( r + w ) * 20 14 = ( r - w ) 30 solve system of equations

OpenStudy (anonymous):

omg thank you so much can you help me with the other problems too please